1.  Lasso Corp. wants to forecast for 2020.  The end of year statements for 2019 are as follows:

Income Statement

Revenues                     $245,622

COGS                            - 142,461

Gross Profits                 103,161

Expenses                       - 49,124

EBIT                                 54,037

- Interest                        -  9,642

EBT                                  44,395

-Taxes                              -7,991

EAT                                  26,637

Balance Sheet

Current Assets           $179,304       Current liabilities          $85,700

Fixed Assets                 $140,930       Long-term debt            78,180

Total Assets                 $320,234       Total liabilities             $163,880

                       Equity                          $156,354

                    Total Liabilities & Equity                      $320,234

Management expects the following for 2020:

An 8% increase in revenues; COGS will increase by 2% from its current percent of sales. Expenses will stay at the same percent of sales;  Current assets will also stay at their same percent of sales. Fixed assets will increase by $18,000 ; All interest will be paid at a 12% interest rate and will only be on long-term debt, which is not expected to change;  the tax rate is expected to stay at the same rate. No dividends are paid and any shortfalls should be made up in current liabilities.

Prepare a forecasted income statement and balance sheet for 2020.  Work to the nearest dollar.
Answer #1

A B с D E F 1 - w N 5 6 7 8 9 10 LASSO CORP. Forecasted Income Statement Year ended 2020 2019 Revenues $245,622 COGS 1424611

H 1 J K L M N 2019 2020 Assets Current Assets Fixed Assets LASSO CORP. Forecasted Balance Sheet Year ended 2020 2020 Liabilit

Notes : 1) All figures taken nearest to dollar.

2) Figures derived for the year 2020 are shown below in formulaes.

А B 1 WN 4 5 6 7 8 LASSO CORP. Forecasted Income Statement Year ended 2020 2019 2020 Revenues 245622 =C6*108/100 COGS 142461

חד G H K 2019 Assets 2019 Current Assets 179304 Fixed Assets 140930 LASSO CORP. Forecasted Balance Sheet Year ended 2020 20203) EAT is Earnings after taxes . But the EAT for 2019 given in the question was $26637 which seems to be not the correct figure. It should be $36404. Because EAT = EBT - Taxes =44395 -7991 =$36404.

4) For any shortfall , it will be adjusted from current liabilities. So Current Liabilities is the balancing figure.

5) Equity contains the EAT and thus adjusted accordingly in 2020.